A RLC circuit is driven by an AC source of 120volt
Answer #1

inductive reactance = \omega L

= 2 x 3.14 x 60 x 100x 10-3

   = 37.68

capacitive reactance = \frac{1}{\omega C}

= 1/ 2 x 3.14 x 60 x 500x 10-6

   = 5.30

Impedence =Z = \sqrt{R^{2}+(X_{c}- X_{L})^{2}}

   = Z = \sqrt{2000^{2}+(1048.5)}

   =2000.26

\omega L = \frac{1}{\omega C}

f = \frac{1}{2\Pi \sqrt{LC}}\

   = 1 / 2 x 3.14 x 7.07 x 10-3

  = 22.52 hz